One of the wagers in the game of roulette is to place a bet that the ball will land on a black number. (Eighteen of the numbers are black, 18 are red, and two are green.) If the ball lands on a black number, the player wins the amount of his bet. If a player bets $4, find the player's expectation. (Round your answer to two decimal places.)
Answer by math_helper(2461)


E = (payoff)*Pr(win) - (wager)*Pr(lose)
= 4*18/(18+18+2) - 4*(18+2)/(18+18+2)
= 72/38 - 80/38
= -8/38

Since this is a negative expectation, the player is expected to LOSE (4/19) dollars or about $0.21 per play.